Determine the best method to solve the system of equations; then solve the system
x=2y-1, 3x-3y=9

Answer by stanbon(75887)

solve the system
x=2y-1, 3x-3y=9
----
Simplify:
x -y = 3
x -2y= -1
---
Subtract and solve for "y":
y = 4
----
y = 4
---
Solve for "x":
x = 2y-1
x = 7
---
Cheers,
Stan H.